OpenStudy (anonymous):

The rectangle has length of 20 inches and width of 80 inches. Find the width in inches of another rectangle of equal area whose length is 40 inches. a. 160 in b. 10 in c. 100 in d. 40 in HELP!!

OpenStudy (anonymous):

20*80=w*40 thusw=20*80/40=40 answer d

OpenStudy (saifoo.khan):

Step 1. The the real area. Step 2. Use that area to find the width.